OpenStudy (anonymous):

(x-1)/5 + 3(x+2) = x/2

OpenStudy (anonymous):

The solution lies in very careful bookkeeping to avoid simple mistakes when simplifying this equation. Let's get started: (X-1)/5 + 3*(X+2) = X/2 Multiply left and righ of the =-sign with 10, to get rid of the '/5' and '/2' divisions: 10(X-1)/5 + 10*3*(X+2) = 10*X/2 Clean up gets us: 2(X-1) + 30*(X+2) = 5*X Now get rid of the brackets; 2*X - 2 + 30*X + 60 = 5*X 32*X + 58 = 5*X Subtract 5*X from both sides; 27*X + 58 = 0, which means: 27*X = -58 Therefore X = -28/27

OpenStudy (anonymous):

sigh...and I made my own mistake at the very bottom. The last line needs to read: Therefore X = -58 / 27 = 2 and 4/27

OpenStudy (anonymous):

thank youu soo soo much, you helped us so much :)